Compare the following 2 alternatives using the Net Present Worth (NPW) method – rate is 4.5% per year. Repeat the solution using the Net Equivalent Uniform Annual (NEUA) method. Which one is simpler? Draw all cash flow diagrams.

Alt. Construction Cost Benefit Maintenance Service Life
A $380,000 $240,000/yr. $10,000/yr. 9 yrs
B $750,000 $270,000/yr. $20,000/yr. 18 yrs
Answer #1

A. Net Present Worth (NPW) method

Present worth analysis will be carried out for 18 yrs (LCM of 9 & 18)

NPW of A = -380000 + (240000 - 10000)*(P/A, 4.5%,18) - 380000*(P/F, 4.5%,9)

= -380000 + 230000*12.159992 - 380000*0.672904

= 2161094.43

NPW of B = -750000 + (270000 - 20000)*(P/A, 4.5%,18)

= -750000 + 250000*12.159992

= 2289997.95

NPW of B is more, so it should be selected

B. Net Equivalent Uniform Annual (NEUA) method

EUAW of A = -380000*(A/P, 4.5%,9) + 240000 - 10000

= -380000*0.137574 + 230000

= 177721.70

EUAW of B = -750000*(A/P, 4.5%,18) + 270000 - 20000

= -750000*0.082237 + 250000

= 188322.32

As EUAW of B is more it should be selected

EUAW is simpler because it does not require LCM analysis and is useful in analyzing projects with unequal life periods

CFD

Alt A

Alt B